OpenStudy (anonymous):

how would I simplify 10x^5y^-3z^8/16x^12y^5z^-2

OpenStudy (anonymous):

........:O, With lots of logs i suppose

OpenStudy (anonymous):

5z^10/8x^7y^8

OpenStudy (anonymous):

10/16=5/8 x^5/x^12=1/x^7 y^-3/y^5=1/y^8 z^8/z^-2=z^10

OpenStudy (anonymous):

u get it??

OpenStudy (anonymous):

yes i do. thanks!